# A representational framework for learning and encoding structurally enriched trajectories in complex agent environments

### v1

@@ Abstract
The ability of artificial intelligence agents to make optimal decisions and generalise them to different domains and tasks is compromised in complex scenarios. One way to address this issue has focused on learning efficient representations of the world and on how the actions of agents affect them, such as disentangled representations that exploit symmetries. Whereas such representations are procedurally efficient, they are based on the compression of low-level state-action transitions, which lack structural richness. To address this problem, we propose to enrich the agent’s ontology and extend the traditional conceptualisation of trajectories to provide a more nuanced view of task execution. Structurally Enriched Trajectories (SETs) extend the encoding of sequences of states and their transitions by incorporating hierarchical relations between objects, interactions and affordances. SETs are built as multi-level graphs, providing a detailed representation of the agent dynamics and a transferable functional abstraction of the task. SETs are integrated into an architecture, Structurally Enriched Trajectory Learning and Encoding (SETLE), that employs a heterogeneous graph-based memory structure of multi-level relational dependencies essential for generalisation. Using reinforcement learning as a data generation tool, we demonstrate that SETLE can support downstream tasks, enabling agents to recognise task-relevant structural patterns across diverse environments.
@@ 1 Introduction
Representation learning plays a crucial role in addressing these challenges by developing structured, transferable encodings of an agent’s environment (Bengio et al.,, 2013; Lesort et al.,, 2018). Various approaches have been proposed, for example, hierarchical RL is particularly effective for hierarchical RL (Pateria et al.,, 2021) introduces structured abstractions for multi-step decision-making but remains constrained by predefined reward structures (Sutton et al.,, 1999; Bacon et al.,, 2017). Goal-conditioned RL enhances planning by leveraging future states as learning targets, yet still depends on explicitly defined objectives (Kaelbling,, 1993; Schaul et al.,, 2015). Disentanglement techniques refine state representations by separating controllable and uncontrollable factors, improving sample efficiency (Higgins et al.,, 2018).
@@
Although these proposals have been instrumental in enhancing policy learning, improving sample efficiency, and facilitating task-specific generalisation, we believe that a new conceptual framework is needed for representational learning that moves beyond compression methods towards enriching structural dependencies between states and actions (Nie et al.,, 2017).
@@
Traditional representations provide a functional foundation, however, they lack the flexibility to support generalisation and knowledge transfer effectively, which may require enriching RL’s ontology. We claim that rather than treating states as undifferentiated wholes, decomposing them into their constituent objects and integrating interaction semantics would yield a more hierarchically structured and transferable representation of task dynamics. This approach would enhance generalisation across tasks, allowing agents to recognise shared structural patterns even in novel environments.
@@
Encoding isolated interactions between objects, however, is not enough; agents need to also learn how these interactions influence future decisions and outcomes. To model this effectively, it is essential to incorporate affordances, which describe the potential interactions available to an agent based on both its capabilities and the environment (Gibson,, 2014). For example, a chair may afford sitting to a human but not to a large animal, such as an elephant. Structured representations of affordances enable an agent to predict the effects of its interactions, providing a foundation for adaptive motor planning and effective task generalisation (Schack and Ritter,, 2013), also enable agents to anticipate how changes in context impact success or failure, reducing the risk of misgeneralisation (Khazatsky et al.,, 2021).
@@
Despite all the advantages that representing affordances would introduce, they alone would only apply to a given state. That is, affordances do not track how interactions unfold sequentially. Tracking how interactions unfold sequentially requires maintaining a structured history of transitions, a trajectory. By retaining this structured trajectory rather than relying solely on the Markov assumption that states encapsulate all relevant history, disregarding transitions (Niv,, 2019), agents will no longer be compelled to make decisions based only on their current state and will be able to learn the dependencies that shape long-term task execution, enabling more adaptive and informed decision-making. We believe that encoding such sequences of interactions and their distant effects is essential for AI systems to adapt to novel scenarios (Eysenbach et al.,, 2022) enabling the agent to capture long-term dependencies, to avoid inefficient learning and goal misgeneralisation (Di Langosco et al.,, 2022).
@@
    A novel architecture SETLE to extract and encode SETs
@@
    Experimental evidence suggesting that SETLE effectively captures both task-specific patterns and cross-task similarities, supporting structured trajectory learning in dynamic, episodic environments  [^1] The focus of this work is creating a representation framework. Evaluating its impact on RL’s policy learning and decision-making is beyond the scope of this work.
The remainder of this paper is structured as follows: first, we review related work on trajectory representation learning and heterogeneous graph models, contextualising the limitations of current approaches. Next, we introduce the SETLE framework, detailing its hierarchical graph construction and encoding. We then present experimental results, demonstrating SETLE’s ability to generalise across tasks. Finally, we discuss the broader implications of our findings and outline future research directions.
@@ 2 Related work
This section critically explores advances in representing the dynamic task structure in complex environments, as well as their focus, successes, and challenges. We present an analysis of outcomes and trajectories highlighting crucial shortfalls and how to address them. Finally, we examine the strengths and limitations of Heterogeneous Graph Neural Networks (HGNNs) in modelling these structured task dynamics.
Most models in representation learning have focused on compression methods which aim to reduce the dimensionality of the data while preserving relevant information from the task, for example transforming high-dimensional data into compact vector representations. In particular, representational learning in agents has been heavily aimed at encoding effective state representations, seeking to capture crucial information about the environment in a compressed and meaningful way (Wang et al.,, 2021). However, while these representations improve efficiency, they do not always generalise well across tasks, as not all aspects of a state can be directly transferred to new environments. In many cases, only certain components of the state remain relevant, while others become task-specific or redundant.
Outcomes, on the other hand, have been considered as a way to summarise the effects of an agent’s actions (Williams,, 1992). Being represented as isolated events, they fail to capture the structured dependencies and hierarchical relationships crucial for generalisation. Trajectories, in contrast, provide a richer representation by encoding the full sequence of states and actions that lead to an outcome while also preserving temporal dependencies. By maintaining the ordering and progression of interactions, trajectories enable agents to understand how decisions influence future states, supporting better generalisation across tasks.
Unlike artificial RL agents, humans do not seem to learn isolated state-action transitions, but rather develop hierarchical models of causal relationships between actions and their effects (Schack and Ritter,, 2013; Morris et al.,, 2022). This enables them to adapt their strategies dynamically, rather than relying on rigid rules or pre-defined goals. Moreover, in real-life scenarios, task dynamics are inherently context-dependent, as the same action can lead to different outcomes depending on environmental conditions (Kiesel and Hoffmann,, 2004).
A key strength of human problem-solving is the ability to use causal models to recognise structural similarities between tasks, transfer knowledge across domains, and apply past experiences to new situations with minimal relearning. For instance, riding a bike and driving a car require distinct motor skills, both involve coordinating movement, adjusting speed, and responding to environmental constraints (Wilson and Collins,, 2019). Representing outcomes plays a crucial role in shaping this learning process. The way in which a task is completed, not just whether it was successful, affects how knowledge is transferred and adapted. If an outcome aligns with prior experiences, it reinforces existing predictive models, while unexpected results prompt adjustments in strategy and refinement of learnt representations.
While outcomes intended or not provide a useful summary of performance, they do not capture the full structure of execution, including the sequential dependencies that lead to a result. Understanding a task requires not just recognising the final state but also interpreting the trajectory of actions taken to reach it, a distinction that is crucial for developing adaptive artificial systems.
A key challenge in reinforcement learning (RL) is learning structured trajectory representations rather than treating experiences as disjointed outcomes. Classically, trajectories in RL are defined as sequences of state-action pairs that evolve over time, capturing the agent’s interaction with the environment from an initial state to termination (Sutton et al.,, 1998). While this definition provides a temporal structure, it does not inherently encode hierarchical relationships beyond stepwise transitions.
One extension of this classical definition is the options framework (Precup,, 2000), which introduces temporally extended actions that allow an agent to execute multi-step behaviours before concluding. An option consists of an internal policy, an initiation set, and a termination condition, enabling an agent to select and commit to more than single-step actions. Other approaches, such as RL with trajectory feedback, shift from evaluating individual steps to assessing the entire trajectory as a whole (Efroni et al.,, 2021). This is particularly useful in scenarios where immediate feedback is sparse or impractical, such as real-world robotics or autonomous systems, where long-term dependencies must be considered. Both options and trajectory feedback offer different ways to extend decision-making beyond stepwise transitions, one by grouping actions into temporally extended choices, the other by evaluating entire trajectories, yet neither provides the level of structural detail necessary for effective generalisation and knowledge transfer. Options, as well as the trajectory feedback approach, treat elements of decision-making as indivisible units, lacking the granularity needed for structured generalisation. Options group actions into fixed temporal sequences, ignoring finer-grained execution details such as object properties, interaction dynamics, and environmental constraints. Similarly, trajectory feedback evaluates entire sequences without analysing individual contributions, making it difficult to identify structural similarities across tasks or extract reusable behavioural patterns that support generalisation and adaptation. This limitation suggests the need for more structured trajectory representations that capture not just individual actions but also high-level task structures and reusable behavioural patterns.
An approach that deals with trajectories in a more structured way is the Self-Consistent Trajectory Autoencoder (SeCTAR) (Co-Reyes et al.,, 2018). SeCTAR frames hierarchical reinforcement learning (HRL) as a trajectory-level representation learning problem. Instead of learning discrete sub-policies, SeCTAR builds continuous latent spaces of trajectories, allowing agents to represent and generate structured behaviours. While SeCTAR’s latent space encodes meaningful trajectory structures, it does not explicitly segment trajectories into transferable skills, potentially limiting its effectiveness for compositional learning and policy transfer. Methods such as the Constructing Skill Trees (CST) algorithm (Konidaris et al.,, 2011) address this limitation by automatically segmenting demonstration trajectories into reusable skills using change-point detection techniques. This approach enables agents to decompose long-horizon tasks into meaningful sub-policies, facilitating skill reuse and hierarchical planning. However, CST relies on human-provided demonstrations, which means that it does not autonomously explore or discover skills without predefined examples. A more generalisable approach would integrate structured trajectory representations with autonomous skill discovery, enabling agents to learn and refine reusable skills without requiring expert demonstrations.
Most existing trajectory representation methods in RL treat trajectories as flat sequences, focussing on linear dependencies rather than hierarchical structures. For example, the Outcome-Driven Actor-Critic (ODAC) framework (Rudner et al.,, 2021), trajectory-based prediction models (Haines et al.,, 2018), and memory-based RL frameworks such as Sequential Episodic Control (SEC) (Freire et al.,, 2024) all adopt flat representations of trajectories as a core design choice. ODAC models latent outcome representations by treating trajectories as flat sequences of state-action transitions, providing a structured yet non-hierarchical approach to inferring task success. Similarly, Haines et al., (2018) emphasises cumulative patterns in state sequences to predict outcomes, assuming linear trajectory structures to simplify trajectory modelling. Meanwhile, SEC improves adaptability and generalisation by retaining episodic state-action pairs, allowing agents to recall past interactions and improve learning efficiency in sparse-reward environments. However, SEC’s flat memory structure prevents it from capturing hierarchical trajectory dependencies, which are essential for structured learning, skill reuse, and transferability.
@@
While these approaches enhance learning by preserving sequential dependencies, they do not inherently encode relational dependencies between entities, their interactions, and their characteristics in a task environment. Understanding how different components interact over time, how specific objects interact or how affordances shape task execution, requires a representation that can capture both temporal and relational dependencies.
Graph-based approaches provide a natural way to represent this structured knowledge, allowing the integration of sequential flow, affordance-driven interactions, and task-specific dependencies into a unified framework.
Heterogeneous Graph Neural Networks (HGNNs) offer a potential solution by using meta-paths, hops, and network schemas to capture relational dependencies between different node types (Wang et al.,, 2021). Unlike traditional graph-based methods, HGNNs can encode multi-level interactions, making them well-suited for tasks that require hierarchical reasoning. Recent work has explored how causal graph representations can further enhance trajectory-oriented learning by explicitly modelling causal dependencies between entities, ensuring that the learnt representations reflect structured transitions rather than mere correlations (Jin et al.,, 2024). Their approach constructs causal chains and applies counterfactual link prediction to refine causal reasoning in structured graphs. However, many HGNN-based causal graph learning approaches, including those in Jin et al., (2024), rely on semi-supervised learning, which assumes access to predefined causal links or labelled annotations. Their effectiveness is often limited in RL environments, where labels such as ground-truth causal effects of actions are typically unavailable or difficult to infer. This makes such methods unsuitable for agents that must learn through interaction, as they require self-supervised or adaptive learning mechanisms.
Heterogeneous Co-contrastive Learning (Wang et al.,, 2021) stands out as a self-supervised HGNN framework that overcomes these limitations by leveraging co-contrastive learning. Unlike traditional approaches, Wang et al., (2021) uses two complementary views: the network schema and the meta-path one, to train node embeddings. By contrasting these views, it captures both local and higher-order structures within the graph, providing a richer representation of static heterogeneous networks without requiring labelled data. Despite its strengths, Wang et al., (2021)’s focus on node-centric representations and static graph structures limits its applicability in dynamic, episodic environments. Such environments require representations that evolve across hierarchical and task-based episodes, reflecting the sequential and cumulative nature of actions leading to outcomes.
To enhance adaptability, Graph Transformer Co-contrastive Learning (GTC) (Sun et al.,, 2025) extends beyond node-centric HGNNs by introducing a dual-branch architecture that integrates local GNN-based aggregation with Transformer-based global modelling, allowing it to capture both fine-grained node relationships and broader contextual dependencies. Additionally, its metapath-aware Hop2Token mechanism encodes multi-hop dependencies more effectively by transforming node interactions into structured token representations, enabling the model to incorporate hierarchical information across varying graph distances. GTC outperforms traditional HGNNs in static settings by capturing high-level semantic information without succumbing to over-smoothing.
Despite recent advancements, existing methods fail to fully address the challenges of dynamic, episodic data, as they do not explicitly capture the hierarchical and cumulative nature of task dynamics. While Haines et al., (2018) and Zou and Suzuki, (2024) each contribute to different aspects of outcome learning, they lack a structured representation that models how trajectories evolve over sequences of interactions. Similarly, trajectory-based models often focus on sequence prediction, but fail to decompose trajectories into their fundamental components, such as objects and their interactions. This lack of decomposition prevents the identification of shared structural patterns across tasks.
Furthermore, graph-based approaches such as Sun et al., (2025) and Wang et al., (2021) have demonstrated the potential of structured representation learning in capturing relational dependencies between entities. However, these methods have not been designed for reinforcement learning and typically aggregate information at a global level, making it difficult to represent task-relevant interactions within individual episodes. Additionally, they do not account for shared entities, such as objects and actions, that could reoccur across different trajectories. This prevents models from recognising common patterns across tasks, limiting their ability to generalise knowledge and adapt to variations in task execution.
These limitations highlight the need to expand the ontology of reinforcement learning to include structured representations of task elements, hierarchical trajectory dependencies, and cross-episode relationships.
The next section introduces SETLE a framework that addresses these gaps by modelling trajectories as structured subgraphs, incorporating objects, actions, and affordances as key components of task execution. We first describe the model’s conceptual ontology and then its implementation.

### v2

@@ Abstract
The ability of artificial intelligence agents to make optimal decisions and generalise them to different domains and tasks is compromised in complex scenarios. One way to address this issue has focused on learning efficient representations of the world and on how the actions of agents affect them in state-action transitions. Whereas such representations are procedurally efficient, they lack structural richness. To address this problem, we propose to enhance the agent’s ontology and extend the traditional conceptualisation of trajectories to provide a more nuanced view of task execution. Structurally Enriched Trajectories (SETs) extend the encoding of sequences of states and their transitions by incorporating hierarchical relations between objects, interactions, and affordances. SETs are built as multi-level graphs, providing a detailed representation of the agent dynamics and a transferable functional abstraction of the task. SETs are integrated into an architecture, Structurally Enriched Trajectory Learning and Encoding (SETLE), that employs a heterogeneous graph-based memory structure of multi-level relational dependencies essential for generalisation. We demonstrate that SETLE can support downstream tasks, enabling agents to recognise task relevant structural patterns across CREATE and MiniGrid environments. Finally, we integrate SETLE with reinforcement learning and show measurable improvements in downstream performance, including breakthrough success rates in complex, sparse-reward tasks.
@@ 1 Introduction
Representation learning plays a crucial role in addressing these challenges by developing structured, transferable encodings of an agent’s environment (Bengio et al.,, 2013; Lesort et al.,, 2018). Various approaches have been proposed, for example, hierarchical RL (Pateria et al.,, 2021) introduces structured abstractions for multi-step decision-making but remains constrained by predefined reward structures (Sutton et al.,, 1999; Bacon et al.,, 2017). Goal-conditioned RL enhances planning by leveraging future states as learning targets, yet still depends on explicitly defined objectives (Kaelbling,, 1993; Schaul et al.,, 2015). Disentanglement techniques refine state representations by separating controllable and uncontrollable factors, improving sample efficiency (Higgins et al.,, 2018).
@@
Although these proposals have been instrumental in enhancing policy learning, improving sample efficiency, and facilitating task-specific generalisation, moving beyond compression methods towards enriching structural dependencies between states and actions is a critical step for more robust and general-purpose intelligence (Nie et al.,, 2017).
@@
Traditional representations provide a functional foundation, however, they lack the flexibility to support generalisation and knowledge transfer effectively, which may require enriching RL’s ontology. Rather than treating states as undifferentiated wholes, decomposing them into their constituent objects and integrating interaction semantics yields a more hierarchically structured and transferable representation of task dynamics. This approach would enhance generalisation across tasks, allowing agents to recognise shared structural patterns even in novel environments.
@@
Encoding isolated interactions between objects, however, is not enough; agents need to also learn how these interactions influence future decisions and outcomes. To model this effectively, it is essential to incorporate affordances, which describe the potential interactions available to an agent based on both its capabilities and the environment (Gibson,, 2014). An affordance is a property of an object in relation to others and from the perspective of an agent that prompts a specific use or interaction. For example, a chair may afford sitting to a human but not to a large animal, such as an elephant. Structured representations of affordances enable an agent to predict the effects of its interactions, providing a foundation for adaptive motor planning and effective task generalisation (Schack and Ritter,, 2013), also enable agents to anticipate how changes in context impact success or failure, reducing the risk of misgeneralisation (Khazatsky et al.,, 2021).
@@
Despite all the advantages that representing affordances would introduce, they alone would only apply to a given state. That is, affordances do not track how interactions unfold sequentially. Tracking how interactions unfold sequentially requires maintaining a structured history of transitions, a trajectory. By retaining this structured trajectory rather than relying solely on the Markov assumption that states encapsulate all relevant history, disregarding transitions (Niv,, 2019), agents will no longer be compelled to make decisions based only on their current state and will be able to learn the dependencies that shape long-term task execution, enabling more adaptive and informed decision-making. Encoding such sequences of interactions and their distant effects has been shown to be essential for AI systems to adapt to novel scenarios (Eysenbach et al.,, 2022), enabling the agent to capture long-term dependencies and avoid inefficient learning or goal misgeneralisation (Di Langosco et al.,, 2022).
@@
    A novel architecture SETLE to extract and encode SETs, including:
@@
    3. (c)
      A comprehensive evaluation demonstrating that integrating SETLE into a reinforcement learning agent leads to substantial improvements in policy learning, sample efficiency, and generalisation.
@@
    Experimental evidence suggesting (in CREATE and MiniGrid) that SETLE effectively captures both task-specific patterns and cross-task similarities, supporting structured trajectory learning in dynamic, episodic environments.
The remainder of this paper is structured as follows: first, we review related work on trajectory representation learning and heterogeneous graph models, contextualising the limitations of current approaches. Next, we introduce the SETLE framework, detailing its hierarchical graph construction and encoding. We then present in Sections 4 and 5 experimental results that first validate the representational power of SETLE’s encoder in CREATE and MiniGrid environments, and second, demonstrate its effectiveness when integrated into a full reinforcement learning loop. Finally, we discuss the broader implications of our findings and outline future research directions.
@@ 2 Related work
A core challenge in reinforcement learning (RL) is generalisation. While agents can master specific tasks, they often struggle to transfer knowledge to new situations. This limitation stems largely from how they represent their experiences. Early approaches focused on learning compressed state representations (Wang et al.,, 2021), aiming to distil high-dimensional observations into compact vectors. However, while efficient, these static snapshots often fail to generalize well, as not all aspects of a state are transferable to new tasks .
A more powerful approach is to learn from trajectories, the full sequence of states and actions leading to a result. Yet, the first and most basic information gleaned from a trajectory is its outcome (Williams,, 1992). Knowing whether a sequence of actions led to success or failure is a crucial learning signal. But relying solely on outcomes is insufficient for transferring knowledge. For instance, the outcome of travelling to a destination can be achieved by either riding a bike or driving a car. While the outcome is the same, the skills required are entirely different (Wilson and Collins,, 2019). Knowing only that both methods were ’successful’ provides no transferable knowledge about how to perform either task. This emphasises that understanding the trajectory, the way in which a task is completed, is crucial for adapting and transferring skills.
Furthermore, even learning from trajectories may not suffice to capture the nuances of information composition and organisation. Classically, they are defined as ”flat” sequences of state-action pairs, a format that lacks deep structural information. This overlooks the rich internal structure that humans intuitively grasp. A key strength of human problem-solving is the ability to develop hierarchical and causal models of our interactions, allowing us to recognize structural similarities and adapt our strategies dynamically (Schack and Ritter,, 2013; Morris et al.,, 2022). The key challenge for AI, therefore, is not just to record trajectories, but to learn to represent their internal structure. This section reviews the evolution of attempts to solve this problem, from early temporal abstractions to modern, heterogeneous graph-based methods.
2.1 Early Trajectory Representations: From Flat Sequences to Temporal Abstractions
Methods like the Outcome-Driven Actor-Critic (ODAC) trajectory-based prediction models, and memory-based RL frameworks such as Sequential Episodic Control (SEC) all adopt simple, shallow sequences of state-action pairs (Rudner et al.,, 2021; Haines et al.,, 2018; Freire et al.,, 2025), focusing on linear dependencies rather than hierarchical structures (Sutton et al.,, 1998).
While these approaches improve learning by preserving sequential patterns, their flat memory structure prevents them from capturing the deeper hierarchical dependencies essential for structured learning and skill reuse.
Early efforts to add structure focused on temporal abstraction. For example, Precup and collaborators’ options framework (Precup et al., 2000) temporally extended actions, allowing agents to execute multi-step behaviours. Similarly, Efroni et al., (2021) work on RL with trajectory feedback, shifted evaluation from single steps to entire trajectories. However, both approaches treated behaviours as indivisible units, lacking the granularity needed for structured generalisation. Fine-grained execution details, including the object properties and interaction dynamics, were disregarded in these approaches, making it difficult to identify reusable behavioural patterns.
2.2 The Shift to Structured Representations: Skills and Latent Spaces
To better capture a trajectory’s internal structure, subsequent research focused on learning representations of entire trajectories. The Self-Consistent Trajectory Autoencoder (SeCTAR) model framed the problem as one of representation learning, building continuous latent spaces of trajectories to represent structured behaviours (Co-Reyes et al.,, 2018). While a significant step, SeCTAR’s latent space does not explicitly segment trajectories into transferable skills, limiting its use for compositional learning.
In contrast, algorithms like Constructing Skill Trees (CST) automatically segment demonstration trajectories into reusable skills (Konidaris et al.,, 2011). CST enabled the decomposition of long tasks into meaningful sub-policies. However, CST’s reliance on human-provided demonstrations means it cannot autonomously discover skills through its own exploration.
2.3 Relational Learning: The Emergence of Graph-Based Approaches
@@
While the aforementioned methods advanced temporal and hierarchical modelling of trajectories, they did not inherently enable the encoding of the relational dependencies between entities within a state. Understanding how different components interact over time requires a representation that can capture both temporal and relational dependencies.
Graph-based approaches provide a natural way to represent this structured knowledge. Heterogeneous Graph Neural Networks (HGNNs) are particularly well-suited, as they can encode multi-level interactions between different types of nodes (e.g., objects, agents, affordances) using mechanisms like meta-paths (Wang et al.,, 2021). Some research has explored causal graphs to model structured transitions rather than mere correlations, though these often rely on semi-supervised learning, making them unsuitable for pure RL settings (Jin et al.,, 2024).
2.4 The State of the Art: Heterogeneous and Self-Supervised Graph Models
Recent advances in HGNNs have focused on self-supervised learning. Heterogeneous Co-contrastive Learning (HeCo) uses two complementary views (network schema and meta-path) to learn rich node embeddings for static graphs (Wang et al.,, 2021). Graph Transformer Co-contrastive Learning (GTC) extends this with a Transformer-based architecture to better capture both local and global dependencies (Sun et al.,, 2025).
Despite their power, these state-of-the-art methods share a critical limitation: they are designed for static, node-centric tasks on a single, large graph. Their focus on node-level representations and fixed graph structures limits their applicability in the dynamic, episodic environments of RL, where representations must evolve to reflect the cumulative nature of actions over a trajectory.
2.5 Identifying the Gap and Positioning SETLE
The evolution of trajectory representations highlights a clear gap in the literature. While trajectory-based models often lack relational detail, advanced graph-based models are not designed for the dynamic, episodic nature of RL, revealing the need for a framework that combines the strengths of both. Existing methods often fail to account for shared entities reoccurring across trajectories, which prevents them from recognising common patterns and limits generalisation.
To overcome these limitations, it is crucial to expand the RL ontology to include structured representations of task elements and cross-episode relationships. SETLE is designed to provide a solution. By modelling individual trajectories as distinct, structured subgraphs with shared components, our framework directly addresses the limitations of prior static, node-centric, or flat-sequence models.